Circular No. 6340

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GK PERSEI M. Ishida, A. Yamashita, H. Ozawa, F. Nagase, and H. Inoue, Institute

of Space and Astronautical Science, communicate on behalf of the ASCA team: “ASCA observed GK Per in outburst during Mar. 4.33–4.87 UT in the energy band 0.5–10 keV, during the period that the optical magnitude was in the range V = 11.8–12.1. The mean counting rates of the SIS and GIS sensors were 1.1 and 1.3 counts/s, respectively. Sinusoidal intensity modulation is clearly seen with a period of 351 s, and the amplitude of the sinusoid is ∼ 30 percent in the band 0.7–10 keV. A periodic intensity variation with a typical time scale of several thousands of seconds is also observed. The phase-averaged spectra of both the SIS and GIS are well represented by a thermal bremsstrahlung continuum with a temperature 18_{−4}^{+6} keV, plus a fluorescent iron emission line centered at 6.40 ± 0.02 keV with an equivalent width of 100 ± 30 eV. No evidence of thermal plasma emission lines is obtained. Both of the spectra show a heavy absorption feature that can be approximated by the sum of three components of different absorptions with hydrogen column densities of 50, 6, and 1 × 10^{22} cm^{−2}. The covering fractions of the three components over the x-ray emission region are 75, 24, and 1 percent, respectively. The x-ray flux at 2–10 keV is 1.9 × 10^{−10} erg cm^{−2} s^{−1}, corresponding to 10 mCrab; assuming a distance of 525 pc to GK Per, we obtain an x-ray luminosity of 6 × 10^{33} erg/s in this energy range. Correcting for the absorption, the bolometric luminosity is estimated to be as large as 3 × 10^{34} erg/s.”

COMET 29P/SCHWASSMANN-WACHMANN 1 H. Boehnhardt and C. Ries, University of Munich; and S. Peschke,

Max-Planck-Institut für Kernphysik, Heidelberg, write: “We performed V and R CCD imaging of comet 29P on Mar. 8–9 with the 0.80-m Mt. Wendelstein telescope of the University of Munich Observatory. The comet showed a large coma of slightly elliptical shape (north-south diameter > 2′.2 = 500 000 km; east-west diameter > 1′.7 = 390 000 km) with a bright central condensation. A strong jet spanning an angle of ∼ 30^{o}–40^{o} was present toward p.a. 45^{o} from the nucleus. The jet curved northward and ended in a cloudy structure ∼ 1′ from the central peak in coma brightness, at p.a. 350^{o}. Some fainter structures were detected in the coma in the southwest quadrant and toward the east. The observed features did not show day-to-day variability (to the level of our on-line analysis).”

1996 March 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green